OpenStudy (anonymous):

A vertical parabola has vertex (1,2) and y-intercept -3. What are its x-intercepts?

OpenStudy (amistre64):

depends on if its open up or down

OpenStudy (amistre64):

if its up; no x ints

OpenStudy (amistre64):

but it opens down cause we got -3 as the y int lol

OpenStudy (anonymous):

vertex is (1,2) and y - intercept 3 looks like \[y=a(x-1)^2 +2\]

OpenStudy (anonymous):

you know that if x = 0 you get -3 so put x = 0 and solve for a

OpenStudy (anonymous):

-3=a(0-1)^2+2\] \[-3=a +2\] \[-5=a\]

OpenStudy (amistre64):

-(5x^2 -10x +3) is what I got

OpenStudy (anonymous):

so your equation is \[y=-5(x-1)^2+2\]

OpenStudy (anonymous):

expand to get probably what amistre (6 to go!) got

OpenStudy (amistre64):

id rather keep it your way and get the sqrts

OpenStudy (anonymous):

actually i think it is \[-5x^2+10x-3\]

OpenStudy (amistre64):

it is, but the negative 1 factors out to make life easier

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ok i wasn't paying attention you are right!

OpenStudy (amistre64):

0=−5(x−1)^2+2 5(x−1)^2 = 2 (x−1)^2 = 2/5 x−1 = +- sqrt(2/5) x = 1 +- sqrt(2/5)

OpenStudy (amistre64):

your equation is quicker to find the xints from :)

OpenStudy (anonymous):

now set = 0 and solve \[-5x^2+10x-3=0\] \[5x^2-10x+3=0\]

OpenStudy (anonymous):

that is for sure. should have left it that way!

OpenStudy (amistre64):

should have lol

OpenStudy (anonymous):

standard form might be \[1\pm \frac{\sqrt{10}}{5}\]

OpenStudy (anonymous):

or even \[\frac{5\pm\sqrt{10}}{5}\]